OpenStudy (anonymous):

-9w *7w^2 = -63w^2 is that right or no check for me and I will medal and fan.

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

-9w *7w^2 = -63w^2 The -63 is correct but not the w^2

OpenStudy (anonymous):

dang

Directrix (directrix):

w^1 times w^2 = ?

OpenStudy (anonymous):

3

Directrix (directrix):

But, what to the third power

OpenStudy (anonymous):

-63w^3

Directrix (directrix):

-9w *7w^2 = -63w^3 Correct

OpenStudy (anonymous):

yay I did it